# Gravitational Wave Signatures of Primordial Black Hole Reheating in Upcoming Interferometry Missions

### v1

@@ 1 Introduction
While the standard reheating scenario assumes that the inflaton field decays into radiation, thereby initiating the radiation-dominated era [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], the PBH reheating scenario presents an alternative, and interesting, possibility. In this framework, PBHs form during a non-standard cosmological epoch characterized by a background equation of state parameter $w_{\phi}$ different from radiation. Depending on the properties of the inflaton potential and the dynamics of the post-inflationary era, the value of $w_{\phi}$ can range between 00 (matter-like domination) and $1$ (stiff fluid domination). In this scenarion, due to their non-relativistic nature, PBHs behave like pressureless matter and can dominate the Universe’s energy density before evaporating via Hawking radiation. This evaporation process injects energy into the cosmic plasma, triggering the reheating of the Universe and marking the onset of the radiation-dominated era [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Such a scenario naturally connects the properties of PBHs to the dynamics of the reheating epoch, offering a novel observational window into the post-inflationary Universe.
A remarkable feature of the PBH reheating scenario is the generation of stochastic GW backgrounds. Two distinct mechanisms contribute to the production of GWs in this context. First, the inhomogeneous spatial distribution of PBHs induces isocurvature density perturbations, which act as secondary sources of GWs during the PBH-dominated era [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. These induced GWs originate from the Poissonian nature of the PBH number density fluctuations. Second, the sudden transition from PBH domination to radiation domination, driven by PBH evaporation, amplifies adiabatic curvature perturbations and generates an additional GW signal [66, 67, 61, 62, 68, 64, 65]. The spectrum of these GWs carries distinctive features that encode information about the PBH parameters, the background equation of state, and the duration of the PBH-dominated era.
@@ 2 Brief overview of the primordial black hole reheating
However, in the scenario we are considering a different background evolution, where PBHs form during the inflaton-dominated phase. We consider this as a formation in some non-standard phase with $w_{\phi}$. Important to note that, after formation, PBHs evolve like non-relativistic matter, causing their energy density to dilute more slowly than the total energy density, since the effective EoS satisfies $0<w_{\phi}\leq 1$  [^1] We exclude exact matter domination ($w_{\phi}=0$) because PBH formation in an early matter-dominated phase would take much longer to form an apparent horizon [74] and preventing PBHs from becoming dominant, what we assume here.. As a result, PBHs could dominate the Universe before decaying, provided their initial abundance is large enough [43, 44, 45, 46, 52]. We thus envision a PBH-dominated era sandwiched between the $w_{\phi}$ dominated phase and the standard radiation-dominated era.
@@
Ultralight PBHs can contribute to the generation of GWs in several ways. For instance, large curvature perturbations which is associated with the PBH formation can induce GWs [91, 92, 93, 94, 95, 96, 97, 98], stochastic GW background can be generated by gravitons produced from the evaporation of PBHs [99], induced GWs through fluctuations in the PBH number density, which are isocurvature in nature [56, 55, 59, 63, 61, 62, 100, 101, 65, 64]. Additionally, the sharp transition from matter-domination (PBH-domination) to radiation-domination, driven by PBH evaporation, enhances induced GWs, which are adiabatic primordial curvature perturbations [102, 66, 61, 62]. Since ultralight PBHs form near the end of inflation, the secondary GWs induced by scalar perturbations from PBH formation typically occur at very high frequencies, close to the wave number associated with end of inflation. The stochastic GW background resulting from PBH evaporation into gravitons also contributes to high-frequency GWs. Thus, we focus on two relevant sources within the detectable range: one related to induced GWs from isocurvature perturbations arising from PBH density fluctuations, and the other from adiabatic fluctuations, which become significant due to the sudden transition from PBH domination to radiation domination. Our ultimate goal is to link the amplitude and frequency dependency of these GWs to PBH and background parameters ($w_{\phi},\beta,M_{\rm in}$), so that any future detection of stochastic GW background can provide insights into both the early PBH domination and the phase preceding it.
@@ 3.1 Signal-to-noise ratio for different detectors
[…]
where $\Omega_{\rm{GW}}(f,\{\theta\})h^{2}$ represents the GW energy density predicted by a specific theory, with $\{\theta\}(\equiv w_{\phi},\beta,M_{\rm in}$ in our case) being the model parameters. $\Omega_{\rm GW}^{\rm noise}(f)h^{2}$ contains noise characteristic of the detector, operating within the frequency-range between $[f_{\rm min},f_{\rm max}]$ and $\tau_{\rm obs}$ denotes the duration of observation. In this analysis, we have considered ET and LISA as the representative missions, for which the noise spectra for instrumental noise are described in App. A. In Table-1, the frequency-range and observation time for the detectors are presented. For both detectors, we have set the observational threshold for the SNR to be $1$.
@@
LISA $\left[10^{-4}-1\right]$ Hz $4$ years
ET $\left[1-10^{4}\right]$ Hz $5$ years
@@ 3.2 Statistical analysis with likelihood function
$\mathscr{L}(\theta)=\prod_{b=1}^{N_{b}}\sqrt{\frac{n_{b}}{2\pi\Omega_{\rm n}(f_{b})^{2}}}\,\,{\rm exp}\left(-\frac{n_{b}\left(\Omega_{\rm sig}(f_{b},\{\theta\})-\Omega(f_{b})\right)^{2}}{\Omega_{\rm n}(f_{b})^{2}}\right).$ (3.3)
where $\Omega_{\rm fid}(f_{b})\equiv\Omega_{\rm sig}(f_{b},\{\theta_{\rm fid})\}$ with $\{\theta_{\rm fid}\}$ indicates the fiducial values of the model parameters. Instead of directly maximizing the likelihood ($\mathscr{L}$), a more convenient approach is to maximize its logarithm, which can be defined as
@@
Although SNR analysis asses the parameter space where is the detection is robust, it does not provide precise measurements of the parameters being probed. Fisher matrix analysis estimates the uncertainties associated with the parameter, offering insight into the precision of the parameter during measurement. The Fisher matrix can be expressed, in terms of the log-likelihood, as [106]
@@
As discussed in the previous section, the Fisher matrix provides insights only into the uncertainties for a given fiducial value. To obtain a more comprehensive inference of the parameters, we perform a rigorous MCMC analysis for the GW detectors ET and LISA. In this context, we employ the same likelihood function as in Eq. (3.3), but instead of $\Omega_{\rm fid}(f_{b})$, we utilize mock data. To generate the mock catalogues, we have chosen a set of parameters such that i) SNR $>1$ and ii) relative uncertainties on the parameters (obtained from Fisher analysis) less than $1$ i.e. $\Delta\theta/\theta<1$. To make it more realistic, we have sampled the chosen set of parameters from a normal distribution, incorporating the noise spectra associated with the detectors [107], ensuring that the mock catalogue is different depending on the spectra and the detectors. The desired posterior distributions of the parameters are expressed as the product of likelihood and priors on the parameters. The prior ranges are tabulated in Table-2. To perform the whole analysis, we have used the publicly available sampler code emcee [108], whereas GetDist [109] is used to plot the posterior distribution.
@@
$\log_{10}(\beta)$ Flat, $-22$rightarrow 00
$\log_{10}(M_{\rm in}/1\,{\rm g})$ Flat, $-0.3$rightarrow $8.62$
@@ 4 Gravitational waves from PBH reheating
[…]
@@ 4.1 Gravitational waves induced by isocurvature fluctuations
$\left\langle\frac{\delta\rho_{\rm BH}(\mathbf{x})}{\rho_{\rm BH}}\frac{\delta\rho_{\rm BH}(\mathbf{\tilde{x}})}{\rho_{\rm BH}}\right\rangle=\frac{4\pi}{3}\left(\frac{d}{a}\right)^{3}\delta(\mathbf{x}-\mathbf{\tilde{x}})\,.$ (4.4)
@@
$\left\langle S_{i}(k)S_{i}(\tilde{k})\right\rangle=\frac{2\pi^{2}}{k^{3}}\,\mathcal{P}_{\rm BH,i}(k)\,\delta(k+\tilde{k}),$ (4.5)
@@ 4.1.1 Statistical analysis in the light of the detectors
Fig. 3 and Fig. 5 depict the SNR in the $w_{\phi}$-$\beta$ plane for different values of $M_{\rm in}$, corresponding to the future GW missions ET and LISA, respectively. The black solid line represents SNR = 1, while the regions where $\rm{SNR}>1$ are indicated by black arrows in both cases. Additionally, the region where $\beta<\beta_{c}$ and the constraints from GW overproduction—derived from the upper bound on $\Delta N_{\rm eff}$ from BBN, with $\Delta N_{\rm eff}=0.28$ based on Planck legacy data [116]—are highlighted in grey and light orange, respectively. In both scenarios, the SNR increases with $\beta$ for a fixed $w_{\phi}$ and vice versa, a trend discussed in detail in the previous section.
@@
On the other hand, for LISA, which is sensitive to frequencies in the range of $0.510^{-5}-$ 0.5 Hz, we find that only PBH masses above $10^{8}$ g are compatible with its detection. However, the SNR values remain very close to 1, even for the maximum allowed PBH formation mass of $\sim 5\times 10^{8}$ g (calculated by considering the evaporation point at the time of BBN), which is also clear from Fig. 5. This SNR-based analysis clearly indicates that, in terms of detection possibilities, the isocurvature source is more favourably detected by ET than LISA.
@@
One of the key findings of the Fisher analysis is that in most of the parameter space where ${\rm SNR}>1$, the relative uncertainties $\Delta w_{\phi}/w_{\phi}$ and $\Delta\beta/\beta$ are both large (i.e., $\Delta w_{\phi}/w_{\phi},\Delta\beta/\beta>10$), as evident in Fig. 6 and Fig. 7. We find that only for masses around $M_{\rm in}=10^{5}$ g, the SNR $>1$ region is compatible with the uncertainty constraint $\Delta w_{\phi}/w_{\phi}<10$, though $\Delta\beta/\beta$ remains greater than 10.
@@ 4.2 Gravitational waves induced by adiabatic fluctuations
During the PBH-domination, the gravitational potential remains constant at the subhorizon scales. After PBH-domination, the gravitational potential oscillates due to radiation pressure in the RD era. During the transition to radiation domination, but before these oscillations begin, the potential on subhorizon scales decays, depending on the transition duration. Since PBH evaporation is nearly instantaneous, the transition is too short for significant decay before oscillations start. As a result, strong GWs are generated following the sudden PBH reheating scenario, even if the primordial curvature perturbation power spectrum remains nearly scale-invariant. We follow Ref. [114] for the analytic expressions, used in our study for this type of GW production. For the adiabatic initial conditions, $\mathcal{P}_{\Phi}$ is related to the initial curvature power spectrum as
@@
with $k_{\rm UV}$ and $k_{\rm IR}$ being the UV and IR cutoff scale, respectively. This changes the Eq. (4.3) as following:
[…]
with $\bar{x}\equiv k\bar{\eta}$ and $\mathcal{I}^{2}_{\rm osc}$ being the oscillatory kernel function, used to estimate the kernel function at the end of PBH-domination, as mentioned in Eq. (4.3). Here the effective spectral index ($n_{\rm eff}$) is related to the scalar spectral index ($n_{s}$) by
@@
For $k\ll k_{\rm BH}$, $n(w_{\phi})=0$ whereas at the other regions, it is given by
with $b\equiv(1-3w_{\phi})/(1+3w_{\phi})$.
For $k\gg k_{\rm BH}$, we can also obtain the resonant contribution for the GW spectrum, which can be written as
@@
In the above expressions $A_{\Phi}(w_{\phi})$ is given by
Here, both $\xi_{1}$ and $\xi_{2}$ are the functions of $w_{\phi}$. We have set $\xi_{1}=1$ throughout our analysis as it is a $\mathcal{O}(1)$ factor [114]. $\xi_{2}$ can be determined as follows:

### v2

@@ 1 Introduction
While the standard reheating scenario assumes that the inflaton field decays into radiation, thereby initiating the radiation-dominated era [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], the PBH reheating scenario presents an alternative, and interesting, possibility. In this framework, PBHs form during a non-standard cosmological epoch characterized by a background equation of state parameter $w_{\phi}$ different from radiation. Depending on the properties of the inflaton potential and the dynamics of the post-inflationary era, the value of $w_{\phi}$ can range between 0 (matter-like domination) and 11 (stiff fluid domination). In this scenarion, due to their non-relativistic nature, PBHs behave like pressureless matter and can dominate the Universe’s energy density before evaporating via Hawking radiation. This evaporation process injects energy into the cosmic plasma, triggering the reheating of the Universe and marking the onset of the radiation-dominated era [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Such a scenario naturally connects the properties of PBHs to the dynamics of the reheating epoch, offering a novel observational window into the post-inflationary Universe.
A remarkable feature of the PBH reheating scenario is the generation of stochastic GW backgrounds. Two distinct mechanisms contribute to the production of GWs in this context. First, on the small scales where the PBHs can be treated as individual particles, the inhomogeneous spatial distribution of PBHs induces isocurvature density perturbations, which act as secondary sources of GWs during the PBH-dominated era [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. These induced GWs originate from the Poissonian nature of the PBH number density fluctuations. On the other hand, on large scales PBHs behave as pressure-less fluid and the sudden transition from PBH domination to radiation domination, driven by PBH evaporation, amplifies adiabatic curvature perturbations and generates an additional GW signal [67, 68, 61, 62, 69, 64, 65]. The distinction between these two sources are purely phenomenological, reflecting differences in scales rather than any physical separation, although both are produced due to sudden changes in the background equation of state from PBH domination to radiation dominated epoch. The spectrum of these GWs carries distinctive features that encode information about the PBH parameters, the background equation of state, and the duration of the PBH-dominated era.
@@ 2 Brief overview of the primordial black hole reheating
However, in the scenario we are considering a different background evolution, where PBHs form during the inflaton-dominated phase. We consider this as a formation in some non-standard phase with a background equation of state parameter $w_{\phi}$. The time-dependent oscillations of the inflaton field can be expressed as
$\phi(t)=\phi_{0}(t)\,\mathcal{P}(t)\,$, (2.1)
where $\phi_{0}(t)$ denotes the oscillation amplitude and $\mathcal{P}(t)$ captures the periodic behaviour. The amplitude $\phi_{0}(t)$ evolves due to redshift and decay, assuming that the oscillation time scale is much shorter than the time scales associated with redshift and decay. Averaging over a single oscillation yields the relation $\langle\dot{\phi}^{2}\rangle\simeq\langle\phi\,dV_{\phi}/d\phi\rangle$, which implies that the inflaton energy density can be approximated as $\rho_{\phi}=\langle\dot{\phi}^{2}/2+V(\phi)\rangle\sim V(\phi_{0})$. Under these assumptions, the averaged equation of state (EoS) of the inflaton is given by [34, 76]
$w_{\phi}=\frac{P_{\phi}}{\rho_{\phi}}=\frac{\langle\dot{\phi}^{2}/2-V(\phi)\rangle}{\langle\dot{\phi}^{2}/2+V(\phi)\rangle}\simeq\frac{n-2}{n+2}\,$, (2.2)
where we assume that the inflaton potential near its minimum takes the form $V(\phi)\sim\phi^{2n}$. For a potential that behaves as $V(\phi)\sim\phi^{6}$ near its minimum, the inflaton exhibits an average background equation of state $w_{\phi}=0.5$. In scenarios where $w_{\phi}<1/3$, we assume that the inflaton decays during the PBH-dominated era that is taken care of by the decay width. On the other hand, when $w_{\phi}>1/3$, the inflaton must likewise remain stable at least until the PBH domination epoch to maintain consistency with the underlying cosmological framework. Important to note that, after formation, PBHs evolve like non-relativistic matter, causing their energy density to dilute more slowly than the total energy density, since the effective EoS satisfies $0<w_{\phi}\leq 1$  [^2] We exclude exact matter domination ($w_{\phi}=0$) because PBH formation in an early matter-dominated phase would take much longer to form an apparent horizon [77] and preventing PBHs from becoming dominant, what we assume here.. As a result, PBHs could dominate the Universe before decaying, provided their initial abundance is large enough [43, 44, 45, 46, 52]. We thus envision a PBH-dominated era sandwiched between the $w_{\phi}$ dominated phase and the standard radiation-dominated era.
@@
Ultralight PBHs can contribute to the generation of GWs in several ways. For instance, large curvature perturbations which is associated with the PBH formation can induce GWs [94, 95, 96, 97, 98, 99, 100, 101], stochastic GW background can be generated by gravitons produced from the evaporation of PBHs [102], induced GWs through fluctuations in the PBH number density, which are isocurvature in nature [56, 55, 59, 63, 61, 62, 103, 104, 65, 64, 66]. Additionally, the sharp transition from matter-domination (PBH-domination) to radiation-domination, driven by PBH evaporation, enhances induced GWs, which are adiabatic primordial curvature perturbations [105, 67, 61, 62]. Since ultralight PBHs form near the end of inflation, the secondary GWs induced by scalar perturbations from PBH formation typically occur at very high frequencies, close to the wave number associated with end of inflation. The stochastic GW background resulting from PBH evaporation into gravitons also contributes to high-frequency GWs. Thus, we focus on two relevant sources within the detectable range: one related to induced GWs from isocurvature perturbations arising from PBH density fluctuations, and the other from adiabatic fluctuations, which become significant due to the sudden transition from PBH domination to radiation domination. Our ultimate goal is to link the amplitude and frequency dependency of these GWs to PBH and background parameters ($w_{\phi},\beta,M_{\rm in}$), so that any future detection of stochastic GW background can provide insights into both the early PBH domination and the phase preceding it.
@@ 3.1 Signal-to-noise ratio for different detectors
$\text{SNR}\;\equiv\;\sqrt{\tau_{\rm obs}\;\int_{f_{\rm min}}^{f_{\rm max}}df\left(\frac{\Omega_{\rm{GW}}(f,\{\theta\})h^{2}}{\Omega_{\rm GW}^{\rm noise}(f)h^{2}}\right)^{2}},$ (3.1)
where $\Omega_{\rm{GW}}(f,\{\theta\})h^{2}$ represents the GW energy density predicted by a specific theory, with $\{\theta\}(\equiv w_{\phi},\beta,M_{\rm in}$ in our case) being the model parameters. $\Omega_{\rm GW}^{\rm noise}(f)h^{2}$ contains noise characteristic of the detector, operating within the frequency-range between $f_{\rm min},f_{\rm max}]$ and $\tau_{\rm obs}$ denotes the duration of observation. In this analysis, we have considered ET and LISA as the representative missions, for which the noise spectra for instrumental noise are described in App. A. In Table-1, the frequency-range and observation time for the detectors are presented. For both detectors, we have set the observational threshold for the SNR to be $1$.
@@
LISA $\left[10^{-4}-1\right]$ Hz 44 years
ET $\left[1-10^{4}\right]$ Hz 55 years
@@ 3.2 Statistical analysis with likelihood function
Although SNR analysis asses the parameter space where is the detection is robust, it does not provide precise measurements of the parameters being probed. Fisher matrix analysis estimates the uncertainties associated with the parameter, offering insight into the precision of the parameter during measurement. We adopted the following form of the Gaussian likelihood function [109] to perform the Fisher analysis:
$\mathscr{L}(\theta)=\prod_{b=1}^{N_{b}}\sqrt{\frac{n_{b}}{2\pi\Omega_{\rm n}(f_{b})^{2}}}\,\,{\rm exp}\left(-\frac{n_{b}\left(\Omega_{\rm sig}(f_{b},\{\theta\})-\Omega_{\rm fid}(f_{b})\right)^{2}}{\Omega_{\rm n}(f_{b})^{2}}\right).$ (3.4)
where $\Omega_{\rm fid}(f_{b})\equiv\Omega_{\rm sig}(f_{b},\{\theta_{\rm fid})\}$ with $\{\theta_{\rm fid}\}$ indicates the fiducial values of the model parameters. In terms of log-likelihood, the Fisher matrix can be expressed as [109]
@@
As discussed in the previous section, the Fisher matrix provides insights only into the uncertainties for a given fiducial value. To obtain a more comprehensive inference of the parameters, we perform a rigorous MCMC analysis for the GW detectors ET and LISA. In this context, we employ the same form of the likelihood function as defined in Eq. (3.4), however, we replace the fiducial model $\Omega_{\rm fid}(f_{b})$ with mock data $\Omega_{\rm mock}(f_{b})$. This substitution is essential in the MCMC framework, where the likelihood function must quantify the agreement between model predictions and (mock) observational data, rather than a predefined fiducial model. To generate the mock catalogues, we have chosen a set of parameters such that i) SNR >1>1 and ii) relative uncertainties on the parameters (obtained from Fisher analysis) less than $1$ i.e. $\Delta\theta/\theta<1$. To make it more realistic, we have sampled the chosen set of parameters from a normal distribution, incorporating the noise spectra associated with the detectors [110], ensuring that the mock catalogue is different depending on the spectra and the detectors. The desired posterior distributions of the parameters are expressed as the product of likelihood and priors on the parameters. The prior ranges are tabulated in Table-2. To perform the whole analysis, we have used the publicly available sampler code emcee [111], whereas GetDist [112] is used to plot the posterior distribution.
@@
$\log_{10}(\beta)$ Flat, 22-22 $\rightarrow$ 0
$\log_{10}(M_{\rm in}/1\,{\rm g})$ Flat, 0.3-0.3 $\rightarrow$ $8.62$
@@ 4 Gravitational waves from PBH reheating
$\overline{\mathcal{P}_{h}(k,\eta)}=8\int_{0}^{\infty}dv\int_{|1-v|}^{1+v}du\left(\frac{\left(1+v^{2}-u^{2}\right)^{2}-4v^{2}}{4uv}\right)^{2}{\cal P}_{\Phi}(uk){\cal P}_{\Phi}(vk)\overline{I^{2}}(x,u,v)\,$, (4.3)
@@ 4.1 Gravitational waves induced by isocurvature fluctuations
[…]
@@
[…]
@@ 4.1.1 Statistical analysis in the light of the detectors
Fig. 3 and Fig. 5 depict the SNR in the $w_{\phi}$-$\beta$ plane for different values of $M_{\rm in}$, corresponding to the future GW missions ET and LISA, respectively. The black solid line represents SNR = 1, while the regions where $\rm{SNR}>1$ are indicated by black arrows in both cases. Additionally, the region where $\beta<\beta_{c}$ and the constraints from GW overproduction—derived from the upper bound on $\Delta N_{\rm eff}$ from BBN, with $\Delta N_{\rm eff}=0.28$ based on Planck legacy data [115], are highlighted in grey and light orange, respectively. In both scenarios, the SNR increases with $\beta$ for a fixed $w_{\phi}$ and vice versa, a trend discussed in detail in the previous section.
@@
On the other hand, for LISA, which is sensitive to frequencies in the range of $10^{-5}-0.5$ Hz, we find that only PBH masses above $10^{8}$ g are compatible with its detection. However, the SNR values remain very close to 1, even for the maximum allowed PBH formation mass of $\sim 5\times 10^{8}$ g (calculated by considering the evaporation point at the time of BBN), which is also clear from Fig. 5. This SNR-based analysis clearly indicates that, in terms of detection possibilities, the isocurvature source is more favourably detected by ET than LISA.
@@
One of the key findings of the Fisher analysis is that in most of the parameter space where ${\rm SNR}>1$, the relative uncertainties $\Delta w_{\phi}/w_{\phi}$ and $\Delta\beta/\beta$ are both large (i.e., $\Delta w_{\phi}/w_{\phi},\Delta\beta/\beta>10$), as evident in Fig. 6 and Fig. 7. We find that only for masses around $M_{\rm in}=10^{5}$ g, the SNR >1>1 region is compatible with the uncertainty constraint $\Delta w_{\phi}/w_{\phi}<10$, though $\Delta\beta/\beta$ remains greater than 10.
@@ 4.2 Gravitational waves induced by adiabatic fluctuations
During the PBH-domination, the gravitational potential remains constant at the subhorizon scales. After PBH-domination, the gravitational potential oscillates due to radiation pressure in the RD era. During the transition to radiation domination, but before these oscillations begin, the potential on subhorizon scales decays, depending on the transition duration. Since PBH evaporation is nearly instantaneous, the transition is too short for significant decay before oscillations start. As a result, strong GWs are generated following the sudden PBH reheating scenario, even if the primordial curvature perturbation power spectrum remains nearly scale-invariant. We follow Ref. [65] in our study for this type of adiabatic induced GWs production. For the adiabatic initial conditions, $\mathcal{P}_{\Phi}$ is related to the initial curvature power spectrum as
@@
with $k_{\rm UV}$ and $k_{\rm IR}$ being the UV and IR cutoff scale, respectively. To avoid complications arising from non-linear evolution, the primordial power spectrum is typically truncated at a high-wavenumber UV cutoff, $k_{\rm UV}$, which excludes modes that become non-linear by the end of the matter-dominated era. Conversely, an infrared (IR) cutoff, $k_{\rm IR}$, is imposed to eliminate large-scale modes that could conflict with CMB observations. These cutoffs ensure that the analysis remains within the regime of linear perturbation theory and observational consistency. To compute the induced GWs from these adiabatic perturbations, we insert the scalar power spectrum $\mathcal{P}_{\Phi_{0}}(k)$ and the transfer function of the adiabatic mode $\Phi(k)$ into the convolution integral that determines the tensor power spectrum (Eq. (4.3)). Unlike the isocurvature case, the adiabatic spectrum depends on both the spectral index $n_{s}$ and the equation-of-state (EoS) parameter $w_{\phi}$, which controls the transfer function. For modes that enter the horizon during the pre PBH domination (stiff or soft) era, the adiabatic transfer function can be approximated by an effective power-law:
$\Phi_{\text{ad}}(k\gg k_{\text{BH}})=A_{\Phi}(w_{\phi})\left(\frac{k}{k_{\rm BH}}\right)^{n(w_{\phi})}\,.$ (4.17)
The effective exponent $n(w_{\phi})$ depends on $w_{\phi}$, given piecewise:
with $b\equiv(1-3w_{\phi})/(1+3w_{\phi})$ and for $k\ll k_{\rm BH}$, $n(w_{\phi})=0$. The amplitude $A_{\Phi}(w_{\phi})$ is determined using a numerical fit given by the following expression
Schematically, this parametrization gives rise to the following integrals over momentum
$\overline{\mathcal{P}_{h,\rm RD}}(k,\eta,\bar{x}\gg 1)\propto$ $\int_{0}^{\infty}dv\int_{|1-v|}^{1+v}du\left(4v^{2}-\left(1+v^{2}-u^{2}\right)^{2}\right)^{2}(uv)^{n_{\rm eff}(w_{\phi})}\,\overline{\mathcal{I}^{2}_{\rm osc}}(\bar{x},u,v)\,$,
with $\bar{x}\equiv k\bar{\eta}$ and $\mathcal{I}^{2}_{\rm osc}$ being the oscillatory kernel function, used to estimate the kernel function at the end of PBH-domination, as mentioned in Eq. (4.3). Here the effective spectral index ($n_{\rm eff}$) is related to the scalar spectral index ($n_{s}$) and the background cosmology via $w_{\phi}$. Combining this with the scalar power spectrum, the effective spectral index in the GW source integral becomes: [65]
@@
This governs the slope of the induced GW spectrum at large $k$, and is negative for $n_{s}\lesssim 1$. For $k\gg k_{\rm BH}$, we obtain the resonant contribution for the GW spectrum, which can be written as
@@
Here, both $\xi_{1}$ and $\xi_{2}$ are the functions of $w_{\phi}$. However, roughly $\xi_{1}$ is a $\mathcal{O}(1)$ parameter [65]. $\xi_{2}$ can be determined as follows: